OpenStudy (anonymous):

Find the value of sqrt 8\9

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow \sqrt{8 \over 9} = {\sqrt{8} \over \sqrt{9}} }\)

OpenStudy (anonymous):

\[\frac{2}{3} \sqrt{2}\]

OpenStudy (lgbasallote):

\[\large \frac{\sqrt 8}{\sqrt 9} \times \frac{\sqrt 9}{\sqrt 9}\]

OpenStudy (anonymous):

72\89?

OpenStudy (lgbasallote):

\(\sqrt 9 \times \sqrt 9 = 9\) \[\sqrt{8 \times 9} = \sqrt{4 \times 2 \times 9} = 2(3)\sqrt 2\]